OpenStudy (anonymous):

how do u describe 16^1/2 in simplify each expression?

OpenStudy (anonymous):

\[16^{1/2} = \sqrt 16 = \sqrt (4^{2}) = 4\]

OpenStudy (anonymous):

\[16^{1/2}-\sqrt{16}=4\]

OpenStudy (anonymous):

how about \[\sqrt{100}?\]

OpenStudy (anonymous):

nick black, your answer should be zero, not 4

OpenStudy (anonymous):

i think nick meant an equal sign instead of a minus, and for MsSmith: \[\sqrt 100 = \sqrt (10^{2} = 10\]

OpenStudy (anonymous):

yeah, well it was one or the other :)

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

np